OpenStudy (sonyalee77):

Which of the following is an equation of the line that passes through the point (-2,3) and has a slope of 1/3

OpenStudy (sonyalee77):

@superdavesuper

OpenStudy (sonyalee77):

Can u show me step by step?

OpenStudy (superdavesuper):

nope....not when @wio is already helping :) thanks @wio!

OpenStudy (anonymous):

Okay first of all, remember what exactly a line is. It has a constant rise over run. 1 / 3 = (y - 3) / (x - (-2))

OpenStudy (anonymous):

Now we have an equation of the line. However this form is not a simplified form, so we have to simplify it.

OpenStudy (sonyalee77):

how?

OpenStudy (anonymous):

Do any of these steps confuse you? Let me know.

OpenStudy (anonymous):

For eqn2 -> eqn3, I multiplied both sides by (x+2)

OpenStudy (the_fizicx99):

You could've just done y-y1=m(x-x1) y-3=1/3(x-(-2)) y-3= 1/3x + 0.66 +3 +3 y = 1/3x + 3.66

OpenStudy (texaschic101):

y = mx + b slope(m) = 1/3 (-2,3) x = -2 and y = 3 now we sub and find b 3 = 1/3(-2) + b 3 = -2/3 + b 3 + 2/3 = b 9/3 + 2/3 = b 11/3 = b so your equation is : y = 1/3x + 11/3

OpenStudy (anonymous):

There is nothing wrong with using the formula m = (y-y_0) / (x-x_0)

OpenStudy (texaschic101):

11/3 is the same as 3.66...just to let you know Sonya

OpenStudy (the_fizicx99):

We basically just gave the same answer in different formats, lol ._.

OpenStudy (texaschic101):

true...I, myself, do not like to use point-slope formula unless asked. But that is just me

OpenStudy (the_fizicx99):

Both ways are easy tbh.

OpenStudy (anonymous):

I like to use the formula that doesn't require you to solve for any variables and then plug them back in. In this case that formula is going to be m = (y-y0)/(x-x0), and y-y0 = m(x-x0) is just a single step after that. It doesn't matter if you do it in the generalized form or after the substitution is done.

OpenStudy (texaschic101):

do it whichever way is easiest for you Sonya :)